Solve the problems below and show the complete solution.

What is the pH of a buffer solution containing 2.0M CH3COOH and 2.0M CH3COONa?
( Ka of CH3CO2H is 1.8 x10ˉ⁵)
1
Expert's answer
2020-12-06T13:54:07-0500

"HA + H_2O \to A^- + H_3O^+"

|HA| = |CH3COOH| = 2.0M

|A-| = |CH3COO-| = 2.0M

Ka = 1.8 x10ˉ⁵

pKa = -logKa = -log(1.8 x10ˉ⁵) = 4.74


From, pH = pKa + log|A-|/|HA|

pH = 4.74 + log2/2 = 4.74 + log1

pH = 4.74 + 0

pH = 4.74